OpenStudy (anonymous):

How do I solve this system of equations by elimination? {-2x + 2y + 3z = 0 {-2x - y + z = -3 {2x + 3y + 3z = 5

OpenStudy (anonymous):

add equation 1 and 3, get \[5y+6z=5\] add equation 2 and 3 get \[2y+4z=5\] and solve those for \(y\) and \(z\)

OpenStudy (anonymous):

How do I solve for them though?

OpenStudy (anonymous):

whoa i made a mistake!

OpenStudy (anonymous):

Thank you so much for helping me! (by the way)

OpenStudy (anonymous):

add 2 and 3 get \[2y+4z=2\] sorry

OpenStudy (anonymous):

no problem, but how do I find the answer? I don't need you to tell me it, I just want to know how to find it.

OpenStudy (anonymous):

\[5y+6z=5\]\[2y+4z=2\] or if you prefer \[5y+6z=5\]\[x+2z=1\]

OpenStudy (anonymous):

I'm following so far

OpenStudy (anonymous):

now you have a choice, you can multiply the second equation by \(-3\) then add to eliminate the \(z\) or you can rewrite the second equation as \[y=1-2z\] the replace in the first equation to get \[5(1-2z)+6z=5\]

OpenStudy (anonymous):

pretty sure you are going to get \(z=0\) then back substitute to find \(y\) and then finally find \(x\)

OpenStudy (anonymous):

actually once you have \(z=0\) it is easy, you know \(y=1-2z\) and since \(z=0\) you get \(y=1\) and so on

OpenStudy (anonymous):

Oh ok, thank you so much!